# Next-Generation Reservoir Computing for Dynamical Inference

### v1

@@ Abstract
We present a simple and scalable implementation of next-generation reservoir computing for modeling dynamical systems from time series data. Our approach uses a pseudorandom nonlinear projection of time-delay embedded input, allowing an arbitrary dimension of the feature space, thus providing a flexible alternative to the polynomial-based projections used in previous next-generation reservoir computing variants. We apply the method to benchmark tasks — including attractor reconstruction and bifurcation diagram estimation — using only partial and noisy observations. We also include an exploratory example of estimating asymptotic oscillation phases. The models remain stable over long rollouts and generalize beyond training data. This framework enables the precise control of system state and is well suited for surrogate modeling and digital twin applications.
@@
$\mathbf{u}_{H}(t)\in\mathbb{R}^{H\times N}$ Measurement vector of history-embedded system states, containing state vectors at times $t,t-\tau,\ldots t-H\cdot\tau$.
@@ I Introduction
Reservoir computing (RC) has emerged as a powerful and versatile machine learning framework for processing temporal data, particularly in tasks involving time series prediction and the inference of latent variables in dynamical systems [3, 4, 5]. As Lu et al. [1] describe, “Reservoir computing is a machine-learning approach that has demonstrated success at a variety of tasks, including time series prediction and inferring unmeasured variables of a dynamical system from measured variables”. In its classical form, RC leverages a high-dimensional, nonlinear dynamical system — referred to as the reservoir — which is driven by an input signal. The reservoir’s response is then post-processed with a linear map to obtain the desired output, commonly referred to as readout. For a general overview of the theory and state of the art, see also Refs. [6, 7].
@@
A key advantage of RC lies in its simplicity: the reservoir itself operates as a static component that does not require training, while only the output layer (readout) is optimized, typically through linear regression. This design allows for implementations of the reservoir in both software — often utilizing recurrent neural networks — and physical systems. Such flexibility has led in a diverse array of physical implementations, including water buckets [8], optical systems [9], electronic delay circuits [10], and photonic devices [11]. The capacity to treat the reservoir as a black box, without needing a detailed model or adjustable internal parameters, sets RC apart from other machine learning paradigms, such as deep learning.
@@
While these systems have demonstrated impressive capabilities [12, 13, 14, 15, 16, 17], classical RC has limitations, particularly in terms of interpretability and control. The internal state of the reservoir is often opaque, and perturbing or probing the system in a controlled manner can be challenging. These limitations have motivated the development of a new paradigm: next-generation reservoir computing (NGRC) [2].
@@
NGRC departs from the traditional architecture by eliminating the explicit reservoir altogether. Instead, it constructs a high-dimensional feature space of dimension, $M$, (corresponding to an implicit reservoir) through a deterministic projection of the input signal. This results in a simple, fully transparent mapping from one time step to the next, where the system state is entirely known and controllable. Letting $u(t)$ denote the input and $r(t)$ the projected state, the evolution becomes a direct function $r(t+1)=F(r(t),u(t))$, enabling precise interventions and measurements. This is particularly advantageous in applications requiring system perturbation and response analysis, such as the computation of phase response curves [18, 19, 20]. See Sec. 1(b) and Fig. 1 for a more detailed comparison of the two methods.
@@
Our work adopts the next-generation reservoir computing framework with an emphasis on scalability and interpretability. In particular, we introduce a pseudorandom nonlinear projection of time-delay embedded input as a simple and dimension-independent alternative to the polynomial-based projections used previously. This projection maintains flexibility across different system sizes while avoiding the need to manage opaque internal reservoir states. The approach is illustrated on benchmark tasks such as attractor reconstruction, bifurcation diagram estimation, and phase recovery, showing stable long rollouts and robustness to partial or noisy data. We further explore asymptotic phase estimation, which to our knowledge has not yet been demonstrated in this context.
@@ II.0.2 Next Generation Reservoir Computing (NGRC)
Next-generation reservoir computing (NGRC) is a simplified alternative to traditional RC. Instead of relying on an explicitly simulated dynamical system as the reservoir, NGRC uses a static, implicit reservoir structure based on nonlinear vector autoregression [2]. The central idea is to take the time-delay embedded state vector,
$\mathbf{u}_{H}(t):=[\mathbf{u}(t),\mathbf{u}(t-\tau),\mathbf{u}(t-2\tau),\dots,\mathbf{u}(t-H\tau)],$
@@
This approach closely aligns with Koopman operator theory [28, 29, 30], which asserts that the dynamics of a nonlinear system can be represented by a linear operator acting on a suitably chosen set of observables. In NGRC, the delay-embedded state vector $\mathbf{u}_{H}(t)$ is mapped into a higher-dimensional space, resulting in the state $\mathbf{r}(t)=P(\mathbf{u}_{H}(t))$, which serves as such a set of observables, effectively lifting the system into a higher-dimensional function space where its evolution becomes approximately linear. Rather than explicitly constructing the Koopman operator, NGRC learns a linear mapping $\mathbf{W}_{\text{out}}$ between $\mathbf{r}(t)$ and the predicted next state $\mathbf{u}(t+\tau)$. The linearity of $\mathbf{W}_{\text{out}}$ is consistent with the Koopman framework, but NGRC sidesteps the need to estimate the full operator. Finally, we note that, while this operation often is referred to as a “projection” in literature, we may more aptly think of this operation as a “lifting”, as it maps the state into a higher-dimensional space rather than projecting onto a subspace.
An advantage of the NGRC architecture is that the time-delay embedding is explicit, eliminating the long transients required in traditional RC to forget intitial conditions. This simplification makes NGRC more practical for manipulating and initializing states, particularly in digital twin settings, where constructing meaningful initial conditions is often intractable due to the “black box” nature of the reservoir.
@@ III.1 Scope of This Work
Using this projection, we train NGRC models on several representative dynamical systems, leveraging both complete and partial observations. We demonstrate that these models accurately reproduce short-term trajectories, long-term attractors, asymptotic phases, and bifurcation diagrams. Notably, we show that bifurcation diagrams can be inferred even when the bifurcation parameter varies in time as a correlated stochastic process — a scenario often arising in non-stationary data collection conditions, similar to the recurrent neural network approach in [31].
Our NGRC models remain stable over long prediction horizons and do not exhibit any transversal instabilities – it seems that even asymptotically the integrator does not drift away and remains on the same attractor [32]. This stability enables us to estimate bifurcation structure directly from a single long trajectory, employing a quasi-continuation approach.
@@ III.2 Projection $P$
In Ref. [2], the projection is achieved by constructing the Volterra series of the measurements $\mathbf{u}_{H}$, a series which contains monomials formed by multiplying the components of the input signal. In Ref. [32], Chebyshev polynomials were used instead. Various processes may thus be employed for this projection, but a limitation of such polynomial-based expansions is that the projection dimension grows with the size of the input vector, so increasing the observation dimension or embedding depth directly increases the number of features. To avoid this dependency, we adopt a pseudorandom projection scheme based on pairwise mixing of input components. This approach makes the feature dimension $M$ a tunable parameter independent of the input size, thereby offering scalability across different system dimensions.
@@
First, we normalize the values of the input vector $\mathbf{u}_{H}$ to lie within the unit interval $(0,1)$. To maintain numerical stability, a small positive margin $\epsilon>0$ is introduced, ensuring the values fall within ($\epsilon,1-\epsilon$) and never leave the unit interval during computation.
@@
$\begin{split}[&u_{1}(t),u_{1}(t-\tau),\ldots,u_{1}(t-H\cdot\tau),\\ &u_{2}(t),u_{2}(t-\tau),\ldots,u_{2}(t-H\cdot\tau),\\ &\quad\quad\quad\vdots\\ &u_{N}(t),u_{N}(t-\tau),\ldots,u_{N}(t-H\cdot\tau)].\end{split}$ (4)
@@
$p_{-N\cdot H},\ p_{-N\cdot H+1},\ \ldots,\ p_{-1},\ p_{0},$
where, for example, $p_{-N\cdot H}=u_{1}(t)$ and $p_{0}=u_{N}(t-H\cdot\tau)$.
@@
where the indices $i$ and $j$ are chosen from a pseudo-random but deterministic sequence, seeded for reproducibility. Any duplicate values are discarded to prevent degeneracy and to ensure that feature matrix reamins full-rank during training.
@@
A notable property of this projection scheme is that it preserves the unit interval: all projected components lie within $(0,1)$, are typically of the same order, and do not aggregate around particular values. Figure 2 shows the distribution of $p_{m}$ values for for a trajectory of the Lorenz system. The values are distributed smoothly and densely across the interval, contrasting monomial-based projections which often produce extreme values, either very small (near zero) or excessively large, depending on the scaling of the input signal.
@@ III.3 Optimizing the linear model
$\mathbf{U}$ $=[\mathbf{u}(t),\ \mathbf{u}(t-\tau),\ \mathbf{u}(t-2\tau),\ \ldots,\ \mathbf{u}(t-T\tau)]$ (6a)
$\mathbf{P}$ $=[P(\mathbf{u}_{H}(t)),\ P(\mathbf{u}_{H}(t-\tau)),\ \ldots,\ P(\mathbf{u}_{H}(t-T\tau))]$ (6b)
@@ III.5 Validation
Additionally, the least squares solution relies on the invertibility of the matrix $\mathbf{P}\mathbf{P}^{T}$, which can become problematic for simple systems or when the projection dimension $M$ is large. A practical way to mitigate this issue is to add a small amount of measurement noise to the original state vectors $\mathbf{u}$, which improves numerical stability and can also act as a form of regularization.
@@
Figure 3: Mean squared error for one-step prediction as a function of the projection dimension $M$, shown for two different measurement noise levels: 0% (top) and 1% (bottom). Without measurement noise (top), the training error decreases initially but diverges from the validation error as $M$ increases, indicating overfitting and numerical instability — particularly beyond $M=300$. Adding 1% Gaussian noise (bottom) helps regularize the fit, leading to training and validation errors that remain comparable across the full range of $M$. This motivates our choice to use 1% noise in all subsequent experiments, as a compromise between stability and predictive accuracy. Too much measurement noise, however, degrades performance, so a trade-off must be considered. All results are computed using long time series of length $T=10^{5}$ for both training and validation.

### v2

@@ Abstract
We present a simple and scalable implementation of next-generation reservoir computing (NGRC) for modeling dynamical systems from time-series data. The method uses a pseudorandom nonlinear projection of time-delay embedded inputs, allowing the feature-space dimension to be chosen independently of the observation size and offering a flexible alternative to polynomial-based NGRC projections. We demonstrate the approach on benchmark tasks, including attractor reconstruction and bifurcation diagram estimation, using partial and noisy measurements. We further show that small amounts of measurement noise during training act as an effective regularizer, improving long-term autonomous stability compared to standard regression alone. Across all tests, the models remain stable over long rollouts and generalize beyond the training data. The framework offers explicit control of system state during prediction, and these properties make NGRC a natural candidate for applications such as surrogate modeling and digital‐twin applications.
@@
$\mathbf{u}_{H}(t)\in\mathbb{R}^{H\times N}$ Measurement vector of history-embedded system states, containing state vectors at times $t,t-\tau,\ldots t-(H-1)\cdot\tau$.
@@ I Introduction
Reservoir computing (RC) has emerged as a powerful and versatile machine learning framework for processing temporal data, particularly in tasks involving time series prediction and the inference of latent variables in dynamical systems [3, 4, 5]. As Lu, Hunt, and Ott [1] describe, “Reservoir computing is a machine-learning approach that has demonstrated success at a variety of tasks, including time series prediction and inferring unmeasured variables of a dynamical system from measured variables”. In its classical form, RC leverages a high-dimensional, nonlinear dynamical system — referred to as the reservoir — which is driven by an input signal. The reservoir’s response is then post-processed with a linear map to obtain the desired output, commonly referred to as readout. For a general overview of the theory and state of the art, see also Refs. [6, 7].
@@
A key advantage of RC lies in its simplicity: the reservoir itself operates as a static component that does not require training, while only the output layer (readout) is optimized, typically through linear regression. This design allows for implementations of the reservoir in both software — often utilizing recurrent neural networks — and physical systems. Such flexibility has led to a wide range of serious physical implementations, including optical systems [8], electronic delay circuits [9], and photonic devices [10]. And to illustrate just how general the reservoir-computing principle is, researchers have even demonstrated it using simple water tanks where surface ripples act as the reservoir [11] — a setup meant more as a proof-of-concept than a practical device. Quantum-inspired implementations of reservoir computing have also been explored in recent work [12]. The capacity to treat the reservoir as a black box, without needing a detailed model or adjustable internal parameters, sets RC apart from other machine learning paradigms, such as deep learning.
@@
While these systems have demonstrated impressive capabilities [13, 14, 15, 16, 17], classical RC has limitations, particularly in terms of interpretability and control. The internal state of the reservoir is often opaque, and perturbing or probing the system in a controlled manner can be challenging. These limitations have motivated the development of a new paradigm: next-generation reservoir computing (NGRC) [2].
@@
NGRC departs from the traditional architecture by eliminating the explicit reservoir altogether. Instead, it constructs a high-dimensional feature space of dimension, $M$, (corresponding to an implicit reservoir) through a deterministic projection of the input signal. This results in a simple, fully transparent mapping from one time step to the next, where the system state is entirely known and controllable. Letting $u(t)$ denote the input and $r(t)$ the projected state, the evolution becomes a direct function $r(t+1)=F(r(t),u(t))$, enabling precise interventions and measurements. This is particularly advantageous in applications requiring system perturbation and response analysis, such as the computation of phase response curves [18, 19, 20, 21, 22]. See Sec. 1(b) and Fig. 1 for a more detailed comparison of the two methods.
@@
Our work adopts the next-generation reservoir computing framework with an emphasis on scalability and interpretability. In particular, we introduce a pseudorandom nonlinear projection of time-delay embedded input as a simple and dimension-independent alternative to the polynomial-based projections used previously. This projection maintains flexibility across different system sizes while avoiding the need to manage opaque internal reservoir states. The approach is illustrated on benchmark tasks such as attractor reconstruction, bifurcation diagram estimation, and phase recovery, showing stable long rollouts and robustness to partial or noisy data. We also include an exploratory demonstration of asymptotic phase estimation using NGRC.
@@ II.0.2 Next Generation Reservoir Computing (NGRC)
Figure 2: Schematic of the NGRC one-step predictor. A normalized, time-delay embedded state $\mathbf{u}_{H}(t)$ is lifted by the nonlinear projection $P$ into a high-dimensional feature vector, which is mapped linearly by $\mathbf{W}_{\text{out}}$ to predict $\mathbf{u}(t+\tau)$. The diagram shows the single scalar signal $\mathbf{u}(t)$ case; for NN observables, $\mathbf{u}_{H}$ has dimension $NH$ and $\mathbf{W}_{\text{out}}$ is an $M\times N$ matrix.
Next-generation reservoir computing (NGRC) is a simplified alternative to traditional RC. A schematic overview of the NGRC one-step prediction mechanism is shown in Fig. 2. Instead of relying on an explicitly simulated dynamical system as the reservoir, NGRC uses a static, implicit reservoir structure based on nonlinear vector autoregression [2]. The central idea is to take the (normalized) time-delay embedded state vector,
$\mathbf{u}_{H}(t):=[\mathbf{u}(t),\mathbf{u}(t-\tau),\mathbf{u}(t-2\tau),\dots,\mathbf{u}(t-(H-1)\tau)],$
@@
This approach closely aligns with Koopman operator theory [30, 31, 32], which asserts that the dynamics of a nonlinear system can be represented by an infinite-dimensional linear operator acting on a suitably chosen set of observables. In NGRC, the delay-embedded state vector $\mathbf{u}_{H}(t)$ is mapped into a higher-dimensional space, resulting in the state $\mathbf{r}(t)=P(\mathbf{u}_{H}(t))$, which serves as such a set of observables, effectively lifting the system into a higher-dimensional function space where its evolution becomes approximately linear. Rather than explicitly constructing the Koopman operator, NGRC learns a linear mapping $\mathbf{W}_{\text{out}}$ between $\mathbf{r}(t)$ and the predicted next state $\mathbf{u}(t+\tau)$. The linearity of $\mathbf{W}_{\text{out}}$ is consistent with the Koopman framework, but NGRC sidesteps the need to estimate the full operator. Finally, we note that, while this operation often is referred to as a “projection” in literature, we may more aptly think of this operation as a “lifting”, as it maps the state into a higher-dimensional space rather than projecting onto a subspace.
An advantage of the NGRC architecture is that the time-delay embedding is explicit, eliminating the long transients required in traditional RC to forget initial conditions. This simplification makes NGRC more practical for manipulating and initializing states, particularly in digital twin settings, where constructing meaningful initial conditions is often intractable due to the “black box” nature of the reservoir.
@@ III.1 Scope of This Work
Using this projection, we train NGRC models on several representative dynamical systems, leveraging both complete and partial observations. We demonstrate that these models accurately reproduce short-term trajectories, long-term attractors, asymptotic phases, and bifurcation diagrams. Notably, we show that bifurcation diagrams can be inferred even when the bifurcation parameter varies in time as a correlated stochastic process — a scenario often arising in non-stationary data collection conditions, similar to the recurrent neural network approach in Ref. [33].
Our NGRC models remain stable over long prediction horizons and do not exhibit any transversal instabilities – it seems that, even asymptotically, the integrator does not drift away and remains on the same attractor. This stability enables us to estimate bifurcation structure directly from a single long trajectory, employing a quasi-continuation approach.
@@ III.2 Projection $P$
In Ref. [2], the projection is achieved by constructing the Volterra series of the measurements $\mathbf{u}_{H}$, a series which contains monomials formed by multiplying the components of the input signal. In Ref. [34], Chebyshev polynomials were used instead. Other NGRC lifting formulations, including kernel-based variants, have also been studied [35, 36]. Various processes may thus be employed for this projection, but a limitation of such polynomial-based expansions is that the projection dimension grows with the size of the input vector, so increasing the observation dimension or embedding depth directly increases the number of features. To avoid this dependency, we adopt a pseudorandom projection scheme based on pairwise mixing of input components. This approach makes the feature dimension $M$ a tunable parameter independent of the input size, thereby offering scalability across different system dimensions.
Our aim in introducing the pseudorandom projection is not to outperform existing feature constructions such as Chebyshev or Volterra expansions. Polynomial-based schemes are known to have near-optimal approximation properties for smooth functions on bounded intervals, so one should not expect a pseudorandom feature map to achieve lower approximation error for a given feature dimension $M$. The motivation for our construction is different: it permits $M$ to be adjusted continuously and independently of the input dimension, without requiring discrete jumps between entire sets of basis functions (e.g. choosing all triplets instead of all pairs sees the number of final features jump drastically from a square number to a cube number). This flexibility is often advantageous in practice, since any loss in approximation efficiency relative to polynomial bases can be offset by modestly increasing $M$.
@@
First, we normalize the values of the input vector $\mathbf{u}_{H}$ to lie within the unit interval $[0,1]$, as illustrated in Fig. 2. To maintain numerical stability, a small positive margin $\epsilon>0$ is introduced, ensuring the values fall within ($\epsilon,1-\epsilon$) and never leave the unit interval during computation.
@@
$\begin{split}[&u_{1}(t),u_{1}(t-\tau),\ldots,u_{1}(t-(H-1)\cdot\tau),\\ &u_{2}(t),u_{2}(t-\tau),\ldots,u_{2}(t-(H-1)\cdot\tau),\\ &\quad\quad\quad\vdots\\ &u_{N}(t),u_{N}(t-\tau),\ldots,u_{N}(t-(H-1)\cdot\tau)].\end{split}$ (4)
@@
$p_{-(N\cdot H-1)},\ p_{-(N\cdot H-2)},\ \ldots,\ p_{-1},\ p_{0},$
where, for example, $p_{-(N\cdot H-1)}=u_{1}(t)$ and $p_{0}=u_{N}(t-(H-1)\cdot\tau)$.
@@
where the indices $i$ and $j$ are chosen from a pseudo-random but deterministic sequence, seeded for reproducibility. Any duplicate values are discarded to prevent degeneracy and to ensure that feature matrix remains full-rank during training.
@@
A notable property of this projection scheme is that it preserves the unit interval: all projected components lie within $[0,1]$, are typically of the same order, and do not aggregate around particular values. Figure 3 shows the distribution of $p_{m}$ values for a trajectory of the Lorenz system. The values are distributed smoothly and densely across the interval, contrasting monomial-based projections which often produce extreme values, either very small (near zero) or excessively large, depending on the scaling of the input signal.
@@ III.3 Optimizing the linear model
$\mathbf{U}$ $=[\mathbf{u}(t+\tau),\ \mathbf{u}(t+2\tau),\ \mathbf{u}(t+3\tau),\ \ldots,\ \mathbf{u}(t+T\tau)]$ (6a)
$\mathbf{P}$ $=[P(\mathbf{u}_{H}(t)),\ P(\mathbf{u}_{H}(t+\tau)),\ \ldots,\ P(\mathbf{u}_{H}(t+(T-1)\tau))]$ (6b)
@@
In practice, it is common to add Tikhonov regularization to improve conditioning and suppress overfitting:
$\mathbf{W}_{\text{out}}=\mathbf{U}\mathbf{P}^{T}\left(\mathbf{P}\mathbf{P}^{T}+\lambda I\right)^{-1},$ (9)
where $\lambda>0$ is a regularization parameter and $I$ the identity matrix.
@@ III.5 Validation
Additionally, the least squares solution relies on the invertibility of the matrix $\mathbf{P}\mathbf{P}^{T}$, which can become problematic for simple systems or when the projection dimension $M$ is large. A practical way to mitigate this issue is to add a small amount of measurement noise to the original state vectors $\mathbf{u}$, which improves numerical stability and can also act as a form of regularization. Conditioning issues and the sensitivity of NGRC regression to the lifted feature representation have been analyzed in Ref. [38], showing how noise and regularization can improve numerical stability.
@@
[…]
Figure 4: Mean-squared one-step prediction error as a function of the projection dimension $M$ under three regularization scenarios: $(a)$ no regularization, $(b)$ Tikhonov regularization with $\lambda=0.01$, and $(c)$ addition of 1% Gaussian measurement noise to the training inputs. $(a)$ Without regularization, errors initially decrease but instabilities appear for larger $M$. $(b)$ Tikhonov regularization suppresses high-$M$ instabilities, but mild overfitting persists, with training errors consistently below validation errors. $(c)$ Adding measurement noise acts as an implicit regularizer by broadening the sampled state space, yielding training and validation errors that remain comparable across the full range of $M$. This motivates the use of 1% measurement noise in subsequent experiments as a practical balance between stability and predictive accuracy. Excessive noise, however, degrades performance, so a trade-off applies. All results use long trajectories ($T=10^{5}$) for both training and validation.
III.6 Long-term stability
While one-step prediction accuracy is informative, it does not guarantee reliable long-term behavior when the predictor is iterated autonomously. Transverse instability has been identified as an important limitation for NGRC-type models in long-term autonomous prediction [34], and recent work has examined both intrinsic stability conditions [39] and complementary post-processing strategies for enhancing long-horizon robustness [40]. Similar stability concerns have also been reported in classical RC, where training noise has been shown to act as a regularizer that improves long-horizon autonomous behavior [41]. We therefore assess long-term stability under three regularization settings: (a) Tikhonov regularization with $\lambda=0.01$ and no measurement noise, (b) the same Tikhonov regularization with $0.1\%$ measurement noise added to the training inputs, and (c) $1\%$ measurement noise without Tikhonov regularization. For each setting we train 100 independent NGRC models and, for each model, generate 10 autonomous prediction trajectories from slightly perturbed initial conditions.
Figure 5 shows representative prediction trajectories together with a reference trajectory from the true ODE system (here we used Lorenz (11)). Because all signals are normalized to lie in $[0,1]$, we consider a trajectory to have diverged once it leaves this interval and terminate its evolution at that point. The fraction of escaped trajectories is plotted alongside the trajectories.
In case (a), most trajectories exhibit rapid transverse instability and leave the admissible range after only a few steps. In case (b), even a very small amount of measurement noise noticeably improves stability, reducing the escape rate and allowing trajectories to remain near the attractor for longer times. In case (c), training with $1\%$ measurement noise yields the most robust behavior: over the time window tested, none of the trajectories escapes the interval, despite the absence of Tikhonov regularization. This suggests that measurement noise acts as an effective implicit regularizer by broadening the sampled state space and suppressing fragile directions. For this reason, we adopt $1\%$ measurement noise in all subsequent experiments.
[…]
Figure 5: Long-term autonomous prediction test for NGRC under three regularization settings: $(a)$ Tikhonov regularization with $\lambda=0.01$ and no measurement noise, $(b)$ the same Tikhonov regularization with $0.1\%$ measurement noise added to the training inputs, and $(c)$ $1\%$ measurement noise without Tikhonov regularization. For each setting, 100 independently trained models are rolled out from 10 slightly perturbed initial conditions, and trajectories are compared with a reference solution of the true ODE system (11). Because all signals are normalized to $[0,1]$, a trajectory is considered divergent once it leaves this interval, at which point the rollout is terminated. The fraction of runs that diverged is shown alongside the trajectories. $(a)$ sees rapid transverse instability and early divergence in most runs; $(b)$ shows substantially reduced escape rates; $(c)$ demonstrates the strongest stability, with no trajectories leaving escaping over the displayed time window.